A true breeding dominant plant (colored, green) is crossed with a true breeding recessive plant (colorless, yellow). The F1’s are then intercrossed to give the following F2 results:

colored,   green (C_G_) 180

colored,   yellow (C_gg) 28

colorless, green (¢¢G_) 20

colorless, yellow (¢¢gg) 188

Perform a Chi-square test to determine whether the genes are linked or unlinked.

What is your X2 value?  
Answer #1

Chi-Squared Tests

Chi-squared test is the statistical method which is used for the determination of the difference between an observed and expected frequency distribution of a true breeding dominant plant (colored, green) is crossed with a true breeding recessive plant (colorless, yellow).

The trait for colored, green (C_G_) is dominant over colorless, yellow (¢¢gg)

Obseved (O)

colored,   green (C_G_) 180

colored,   yellow (C_gg) 28

colorless, green (¢¢G_) 20

colorless, yellow (¢¢gg) 188

Expected (E)

colored,   green (C_G_) 234 (416x9/16)

colored,   yellow (C_gg) 78 (416x3/16)

colorless, green (¢¢G_) 78 (416x3/16)

colorless, yellow (¢¢gg) 26 (416x1/16)

Now we can calculate the chi-squared test by the following formula:

X2=\sum(O-E)2/E

Now we can calculat O-E for each-

colored,   green (C_G_) (-54)

colored,   yellow (C_gg) (-50)

colorless, green (¢¢G_) (-58)

colorless, yellow (¢¢gg) (162)

Now we can calculate (O-E)2/E for each

colored,   green (C_G_) (12.46)

colored,   yellow (C_gg) (32.05)

colorless, green (¢¢G_) (43.12)

colorless, yellow (¢¢gg) (1009)

X2= (12.46 + 32.5 + 43.12 + 1009) = 1096.63

A chi-squared test require 2 hypotheses:

Null hypothesis (H0): If there is no significany in between the observed and expected frequencies of the trait then the genes are said to be unlinked.

§ Alternative hypothesis (H1): If there is a significany in between the observed and expected frequencies of the trait then the genes are said to be linked.